# Md Kowsher School of Engineering and Science

### v1

@@ 6 Conclusion
We introduce nonlinear readout to reservoir computing that significantly enhances the performance of the state-of-the-art DNNs, including Transformer. Our deep Reservoir Transformer has the unique advantage of handling any input length and producing robust prediction outputs. Thus, it is ideal for chaotic time-series forecasting. This research will open new research questions on seamlessly integrating chaotic analysis, DNNs, and reservoir computing to better understand and predict our environments.
Appendix A Appendix
A.1 Dataset and Parameter Configuration
@@
Dataset Partitioning: The dataset is divided sequentially into three distinct sets: a training set, a validation set, and a test set. Specifically, the first $70\%$ of the data is assigned to the training set, while $10\%$ is allocated to the validation set, and the remaining $20\%$ forms the test set. This partitioning strategy enables effective model training, validation, and evaluation.
Model Parameters: The configuration of model parameters plays a crucial role in our research methodology. For the Transformer architecture, we employ four Transformer blocks, each comprising 32 multi-head attention mechanisms with a total of four attentions. In the case of the Feedforward Neural Network (FNN), we utilize hidden layers containing 64 units each.
@@
Furthermore, the reservoir state size varies between 20 and 50 units across different ensembling reservoirs. To introduce diversity in the ensembling process, sparsity levels range from 0.3 to 0.7 for different reservoirs, accompanied by leaky units with values ranging from 0.2 to 0.6.
Finally, a learning rate of $1\times 10^{-4}$ is adopted to facilitate the training process and achieve optimal convergence.
@@
  • Absenteeism at Work:  [^5] https://archive.ics.uci.edu/ml/datasets/Absenteeism+at+work The dataset focusing on absenteeism at work is sourced from the UCI repository, specifically the ”Absenteeism at work” dataset. In our analysis, the dataset has been utilized for classification purposes, specifically targeting the classification of individuals as social drinkers. The dataset comprises 21 attributes and encompasses a total of 740 records. These records pertain to instances where individuals were absent from work, and the data collection spans the period from January 2008 to December 2016.
@@
A.2 Pseudo Algorithm For Deep Reservoir Transformer
@@
$\vec{u}(1:T,1:c),\vec{y}(1:T)$
Optimize $Pr(\vec{y}(t+1)|\vec{u}(1:t+1),\vec{y}(1:t))$ distribution by learning a model $M(.)$
$\vec{W}_{in},\hat{\vec{W}}\leftarrow\mathcal{N}(0,1)$triangleright Weights initialization of Reservoir
$\sigma\leftarrow\mathcal{N}(0,1)$triangleright Weights initialization of Reservoir non-linear output
$\phi\leftarrow\mathcal{N}(0,1)$triangleright Weights initialization of Transformer encoder $M(.)$
@@
     for $i\leftarrow 1$ to $T$ do
         for $k\leftarrow 1$ to $GroupESN$ do
              $o^{k}(i)\leftarrow R_{k}(\vec{u}(i),\vec{y}(i);\vec{W}_{in},\hat{\vec{W}},\sigma)$triangleright Non-linear readout from $k-th$ reservoir
              $o(i)\leftarrow o(i)\oplus o^{k}(i)$triangleright Element-wise addition among all reservoirs
@@
         $\vec{z}(i+1)\leftarrow\kappa\vec{u}(i+1)+(1-\kappa)o(i+1)$triangleright Concatenation
         $\bar{\vec{y}}(i+1)\leftarrow\mathcal{M}(\vec{z}(i+1);\phi)$
     end for
     $loss\leftarrow\mathcal{L}(\bar{\vec{y}}(i+1),\vec{y}(i+1))$triangleright Calculating loss for every time step $z$
     $\phi,\sigma\leftarrow update$triangleright Update parameters
@@
The proposed algorithm 1 is for training a deep reservoir computing model, with the goal of optimizing the distribution of the target variable, $\vec{y}(t+1)$, given the input variables, $\vec{u}(1:t+1)$, and previous target variables, $\vec{y}(1:t)$. The algorithm initializes the weights of the reservoir, $\vec{W}_{in}$ and $\hat{W}$, as well as the weights of the non-linear output, $\sigma$, and the Transformer encoder, $\phi$, as random variables sampled from a normal distribution with mean 0 and standard deviation 1.
The algorithm then enters an outer loop, where the number of iterations is determined by the number of epochs. Within this loop, there is an inner loop that iterates through each time step, $i$, from 1 to $T$. For each time step, a group of reservoirs, $GroupESN$, are processed in parallel. Each reservoir, $R_{k}$, receives the input, $\vec{u}(i)$, and the previous target variable, $\vec{y}(i)$, as input, and calculates a non-linear readout, $o^{k}(i)$, using the weights $\vec{W}_{in},\hat{\vec{W}}$, and $\sigma$. The readouts from all reservoirs are then element-wise added together to produce $o(i)$.
The input of the next time step, $\vec{u}(i+1)$, is concatenated with the readout from the reservoirs, $o(i)$, using a parameter $\kappa$, to produce $\vec{z}(i+1)$. Finally, the Transformer encoder, $\mathcal{M}(.)$, is applied to $\vec{z}(i+1)$ using the parameters $\phi$, to produce the target variable for the next time step, $\vec{y}(1+1)$.
The loss is calculated for every time step, $z$, by comparing the predicted target variable, $\bar{\vec{y}}(1:T)$, with the true target variable, $\vec{y}(1:T)$, using a loss function, $\mathcal{L}(.)$. The parameters, $\phi$ and $\sigma$, are then updated to minimize the loss.
A.3 Variable Descriptions
@@
$\mathcal{M}(.)$
A Transformer or DNN model.
$\vec{u}$
Samples with $c$ components.
$c$
Number of features in dataset
$\vec{y}$
Target features for prediction.
$\bar{\vec{y}}$
Prediction value
$T$
Entire time span or duration
$t$
Represents a specific moment or data point in time
$R$
Non-linear readout representation of reservoir
$\vec{o}$
Aggregate output from all reservoirs within the group
$\rho$
Spectral radius of reservoir
$Q$
Number of reservoirs in GroupESN
$m$
Reservoir output size
$\vec{W}_{i}n$
Input-to-reservoir weight matrix
${\vec{W}_{out}}$
Reservoir to readout weight matrix
${\theta_{out}}$
Reservoir to readout bias matrix
$\vec{\hat{W}}$
Recurrent reservoir weight matrix
$\vec{x}$
Reservoir state
$\kappa$
Parameter to indicate the weights of current state and reservoir state
$\phi$
Learnable parameters of $M(.)$
$s$
Window size to look back previous sequence
$R(.)$
Deep reservoir computing
$d$
Feature dimension of Transformer
$\vec{g}$
Input for the baseline Transformer
$\vec{z}$
The concatenation of reservoir and current multivariate
$\mathcal{L}$
Loss function
$D_{2}$
Correlation dimension
$C(\epsilon)$
Correlation sum
$H(.)$
Heaviside step function
@@
A.4 Ablation Study
Figure 6: The graph illustrates the relationship between the number of ensemble reservoirs used and the mean square error (MSE) of the prediction outputs. The X-axis represents the number of reservoirs included in the ensemble, while the Y-axis depicts the corresponding MSE values. The ’star’ in the graph denotes the minimum MSE achieved.
A.4.1 Impact of Reservoir Count on Performance:
Figure 6 illustrates the relationship between the number of ensembling reservoirs and system performance. The x-axis depicts the count of reservoirs, while the y-axis portrays the Mean Squared Error (MSE) value. The graph discernibly demonstrates that increasing the number of reservoirs positively correlates with performance enhancement, leading to a reduction in loss. However, beyond a certain threshold of reservoir count (best hyperparameters of the number of reservoirs), the performance improvement saturates, resulting in a plateau in loss reduction. This phenomenon suggests an optimal range for reservoir count in relation to performance enhancement.
A.4.2 Graphical Explanation of Features
In Figure 7, we present a visual depiction that offers a comprehensive insight into the interpretability of features derived from both the reservoir output $\vec{o}(t)$ and input feature $\vec{u}(t)$.
A.5 Experiments on Univariate Time sSeries
@@
Table 7 shows that RT model outperforms GRIN (Cini et al., 2021), BRITS (Cao et al., 2018), ST-MVL (Yi et al., 2016), M-RNN (Yoon et al., 2018), ImputeTS (Moritz & Bartz-Beielstein, 2017) on the air quality dataset. On regression and classification tasks, RT consistently outperforms the baseline Transformer with respect to all criteria, achieving significantly lower MSE and MAE values across various datasets like DWV, DGP, DDFO, and BTC. This suggests that the RT model is better equipped to capture complex temporal patterns. In the classification tasks, the RT model demonstrates better performance in terms of accuracy, F1 Score, Jaccard Score, Roc Auc, Precision, and Recall across different datasets like Absenteeism at work and Temperature Readings. Through this analysis, we provide evidence that the RT architecture excels not only in regression tasks but also in classification tasks, outperforming traditional Transformer architectures.
@@
Table 7: Assessing the performance of regression on both training and test datasets, we utilize a window size of $s=5$ to observe historical data in baseline Transformers.

### v2

@@ 6 Conclusion
Our RT has the unique advantage of handling any input length and producing robust prediction outputs. Thus, it is ideal for time-series forecasting, including event prediction in chaos. We introduce nonlinear readouts to our Reservoir Transformer that outperform state-of-the-art models in terms of accuracy, keeping reasonable complexity in space and time. Furthermore, we testify to long-horizon forecasting and show that RT makes the far-horizon prediction less sensitive to initialization. This research direction has the potential for succession works by integrating chaotic analysis, deep neural networks, and reservoir computing to understand and predict our environments better.
Appendix A Dataset and Parameter Configuration
@@
Dataset Partitioning: For the datasets ETTh1, ETTh2, ETTm1, ETTm2, Traffic, Weather, and ILI, we adopted the data preparation settings outlined by [Zeng et al., 2023]. For other datasets, we divided the data sequentially into three distinct sets: training (70%), validation (10%), and testing (20%). This division strategy facilitates efficient training, validation, and testing of models.
Tools and Hardware: We have used Pytorch and Huggingface libraries to implement our framework. All the experiments were carried out on a single Nvidia RTX-4090 or Tesla H100 GPU.
Model Parameters: The configuration of model parameters plays a crucial role in experiments. For the Transformer architecture, we employ four Transformer blocks, each comprising 12 multi-head attention mechanisms with a total of 12 attention layers. In the case of the Feedforward Neural Network (FNN), we utilize hidden layers containing 64 units each.
@@
Furthermore, the reservoir state size varies between 200 and 300 units across different ensembling reservoirs. To introduce diversity in the ensembling process, spectral radius values range from 0.5 to 0.9 for different reservoirs, accompanied by leaky units with values ranging from 0.2 to 0.6.
Finally, a learning rate of $1\times 10^{-5}$ is adopted to facilitate the training process and achieve optimal convergence.
@@
  • Absenteeism at Work:  [^5] https://archive.ics.uci.edu/ml/datasets/Absenteeism+at+work The dataset focusing on absenteeism at work is sourced from the UCI repository, specifically the "Absenteeism at work" dataset. In our analysis, the dataset has been utilized for classification purposes, specifically targeting the classification of individuals as social drinkers. The dataset comprises 21 attributes and encompasses a total of 740 records. These records pertain to instances where individuals were absent from work, and the data collection spans the period from January 2008 to December 2016.
@@
Appendix B Algorithm for RT
@@
$\mathbf{u}_{1:T}$, a Transformer model $\mathbb{M}(.)$, a groups reservoirs $\mathcal{R}(.)$, an embedding function $\epsilon(.)$
Initialization $\mathbb{M}(.)$ and $\mathcal{R}(.)$ with parameters
@@
     for $t\leftarrow k$ to $T$ do
         $\mathbf{h}_{t}=\epsilon(\mathbf{u}_{t-k:t})$triangleright Embedding the time input(3.1)
         $\mathbf{z}_{t}=\mathcal{R}(\mathbf{h}_{t})$triangleright Readout from Group reservoir (3.2, 3.3, 3.4, 3.5, 3.6 )
         $\mathbf{f}_{t}=(1-\kappa)\mathbf{z}_{t}\uplus\kappa\mathbf{h}_{t}$triangleright Concatenation (3.7)
         $\mathbf{\hat{u}}_{t+1:t+\tau}\leftarrow\mathbb{M}(\mathbf{f}_{t})$
@@
     $loss\leftarrow\mathcal{L}(\mathbf{u}_{t+1:t+\tau},\mathbf{\hat{u}}_{t+1:t+\tau})$triangleright Calculating loss (3.8)
     update parameters
@@
In the pseudo algorithm 1, we describe how to train RT models. The algorithm initiates by setting parameters for both the Transformer and the reservoir groups. During each training epoch, it processes sequences incrementally from a specified start time $k$ up to the total length $T$. Each sequence segment transforms an embedded vector, which is then input into the reservoirs to produce intermediate representations. These are subsequently merged with the original embedded input, controlled by a weighting parameter $\kappa$, forming a composite feature vector. This vector is fed into the Transformer to predict future sequence values. The training cycle completes by calculating loss between the predicted and actual values and updating the model parameters accordingly, aiming to minimize prediction errors in following epochs.
Appendix C Variable Descriptions
We have described the used variables and math notations in Table 9.
@@ Appendix C Variable Descriptions
$\mathbb{M}(.)$ Transformer model
$\mathcal{R}(.)$ Deep reservoir computing
$\mathbf{u}_{t}$ Input at $t$ time steps
$\mathbf{x}_{t}$ Reservoir state at $t$ time steps
$\mathbf{y}_{t}$ Linear output of reservoir
$\mathbf{h}_{t}$ Embedding of time steps
$\mathbf{\Lambda}_{t}$ Non-linear readout
$\mathbf{z}_{t}$ Self attention of group reservoir
$\mathbf{o}_{t}$ Ensembling of group reservoir
$\mathbf{f}_{t}$ Input of Transformer
$T$ Total time input length
$k$ Look back window
$H$ Number of attention heads
$N_{r}$ Reservoir size
$N_{u}$ Reservoir input size
$m$ Reservoir output size
$L$ Number of reservoirs in Group
$h$ Forecasting horizons
$d_{\epsilon}$ Model dimension
$d_{key}$ Dimension of key in attention
$d_{value}$ Dimension of key in attention
$l_{y}$ Number of layers in Transformer
$c$ Compressive Transformer memory size
$r$ Compression ratio
$p$ Number of soft-prompt summary vectors
$s$ Kernel size
$P$ Patch size
$\mathbf{\Phi}$ Transformer’s learnable parameters
$\mathbf{W}_{q}$ Queries weights
$\mathbf{W}_{v}$ Values weights
$\mathbf{W}_{k}$ Keys weights
$\mathbf{W}_{in}$ Input-to-reservoir weight matrix
$\mathbf{W}$ Reservoir weight matrix
$\boldsymbol{\theta}$ Bias-to-reservoir weight
$\mathbf{W}_{out}$ Reservoir-to-readout weight matrix
$\boldsymbol{\theta}_{out}$ Bias-to-readout weight
@@
Appendix D Baselines
We show our experimental results of Reservoir Transformer (RT) compared to baselines, including state-of-the-art methods in time-series prediction, such as FEDformer [Zhou et al., 2022], Autoformer [Wu et al., 2021], Informer [Zhou et al., 2021], Pyraformer [Liu et al., 2021], log(T)rans [Li et al., 2019], GRIN [Cini et al., 2021], BRITS [Cao et al., 2018], STMVL [Yi et al., 2016], M-RNN [Yoon et al., 2018], ImputeTS [Moritz and Bartz-Beielstein, 2017], and Transformer Vaswani et al. [2017]. To conduct experiments on multivariate and uni-variate time series, we evaluate methods outlined by  [Zeng et al., 2022], which includes data splitting, pre-processing, and parameter setting.
Appendix E Long Uni-Variate Forecasting
In Table 10 we have compared the long uni-variate time-series forecasting with state of the art time series method and showed improvement of RT in uni-variate forecasting.
Methods RT Linear NLinear DLinear FEDFormer-f FEDFormer-w Autoformer Informer log(T)rans
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.050 0.172 0.189 0.359 0.053 0.177 0.056 0.180 0.079 0.215 0.080 0.214 0.071 0.206 0.193 0.377 0.283 0.468
192 0.064 0.201 0.078 0.212 0.069 0.204 0.071 0.204 0.104 0.245 0.105 0.256 0.114 0.262 0.217 0.395 0.234 0.409
336 0.081 0.225 0.091 0.237 0.081 0.226 0.098 0.244 0.119 0.270 0.120 0.269 0.107 0.258 0.202 0.381 0.386 0.546
720 0.097 0.292 0.172 0.340 0.080 0.226 0.189 0.359 0.142 0.299 0.127 0.280 0.126 0.283 0.183 0.355 0.475 0.629
ETTh2 96 0.127 0.275 0.133 0.283 0.129 0.278 0.131 0.279 0.128 0.271 0.156 0.306 0.153 0.306 0.213 0.373 0.217 0.379
192 0.168 0.324 0.176 0.330 0.169 0.324 0.176 0.329 0.185 0.330 0.238 0.380 0.204 0.351 0.227 0.387 0.281 0.429
336 0.192 0.351 0.213 0.371 0.194 0.355 0.209 0.367 0.231 0.378 0.271 0.412 0.246 0.389 0.242 0.401 0.293 0.437
720 0.369 0.411 0.292 0.440 0.225 0.381 0.276 0.426 0.278 0.420 0.288 0.438 0.268 0.409 0.291 0.439 0.218 0.387
ETTm1 96 0.024 0.119 0.028 0.125 0.026 0.122 0.028 0.123 0.033 0.140 0.036 0.149 0.056 0.183 0.109 0.277 0.049 0.171
192 0.037 0.149 0.043 0.154 0.039 0.149 0.045 0.156 0.058 0.186 0.069 0.206 0.081 0.216 0.151 0.310 0.157 0.317
336 0.050 0.168 0.059 0.180 0.052 0.172 0.061 0.182 0.084 0.231 0.071 0.209 0.076 0.218 0.427 0.591 0.289 0.459
720 0.072 0.199 0.080 0.211 0.073 0.207 0.080 0.210 0.102 0.250 0.105 0.248 0.110 0.267 0.438 0.586 0.430 0.579
ETTm2 96 0.062 0.182 0.066 0.189 0.063 0.182 0.063 0.183 0.067 0.198 0.063 0.189 0.065 0.189 0.088 0.225 0.075 0.208
192 0.092 0.224 0.094 0.230 0.090 0.223 0.092 0.227 0.102 0.245 0.110 0.252 0.118 0.256 0.132 0.283 0.129 0.275
336 0.115 0.258 0.120 0.263 0.117 0.259 0.119 0.261 0.130 0.279 0.147 0.301 0.154 0.305 0.180 0.336 0.154 0.302
720 0.356 0.382 0.175 0.320 0.170 0.318 0.175 0.115 0.178 0.325 0.219 0.368 0.182 0.335 0.300 0.435 0.160 0.321
Table 10: Univariate long-term forecasting results
Appendix F More results
@@
Table 11 shows that RT model outperforms GRIN [Cini et al., 2021], BRITS [Cao et al., 2018], ST-MVL [Yi et al., 2016], M-RNN [Yoon et al., 2018], ImputeTS [Moritz and Bartz-Beielstein, 2017] on the air quality dataset. On regression and classification tasks, RT consistently outperforms the baseline Transformer with respect to all criteria, achieving significantly lower MSE and MAE values across various datasets like DWV, DGP, DDFO, and BTC. This suggests that the RT model is better equipped to capture complex temporal patterns. In the classification tasks, the RT model demonstrates better performance in terms of accuracy, F1 Score, Jaccard Score, Roc Auc, Precision, and Recall across different datasets like Absenteeism at work and Temperature Readings. Through this analysis, we provide evidence that the RT architecture excels not only in regression tasks but also in classification tasks, outperforming traditional Transformer architectures. Additionally, we have shown the comparison of multi-variate time series with other methods such as LightTS [Campos et al., 2023] and Autformer [Wu et al., 2021] in Table 11.
@@
Table 11: Assessing the performance of regression on both training and test datasets, we utilize a window size of $100$ to observe historical data in baseline Transformers and RT.
Methods
RT
DLinear
PatchTST
TimesNet
FEDformer
Stationary
ETSformer
LightTS
Autoformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.362 0.392 0.376 0.397 0.375 0.399 0.370 0.399 0.384 0.402 0.376 0.419 0.449 0.479 0.424 0.432 0.449 0.459
192 0.396 0.412 0.405 0.416 0.413 0.421 0.436 0.429 0.420 0.448 0.534 0.504 0.538 0.504 0.475 0.462 0.500 0.482
336 0.427 0.422 0.439 0.443 0.422 0.436 0.491 0.469 0.459 0.465 0.588 0.535 0.574 0.521 0.518 0.488 0.521 0.496
720 0.441 0.455 0.472 0.490 0.447 0.466 0.521 0.500 0.506 0.507 0.643 0.616 0.562 0.535 0.547 0.533 0.514 0.512
avg 0.406 0.418 0.422 0.437 0.413 0.430 0.458 0.450 0.440 0.460 0.570 0.537 0.542 0.510 0.491 0.479 0.496 0.487
ETTh2 96 0.262 0.320 0.289 0.353 0.274 0.336 0.340 0.374 0.358 0.397 0.476 0.458 0.340 0.391 0.397 0.437 0.346 0.388
192 0.331 0.373 0.383 0.418 0.339 0.379 0.402 0.414 0.429 0.439 0.512 0.493 0.430 0.439 0.520 0.504 0.456 0.452
336 0.358 0.403 0.448 0.465 0.329 0.380 0.452 0.452 0.496 0.487 0.552 0.551 0.485 0.479 0.626 0.559 0.482 0.486
720 0.369 0.411 0.605 0.551 0.379 0.422 0.462 0.468 0.463 0.474 0.562 0.560 0.500 0.497 0.863 0.672 0.515 0.511
avg 0.330 0.376 0.431 0.446 0.330 0.379 0.414 0.427 0.437 0.449 0.526 0.516 0.439 0.452 0.602 0.543 0.450 0.459
ETTm1 96 0.272 0.336 0.299 0.343 0.290 0.342 0.338 0.375 0.379 0.419 0.386 0.398 0.375 0.398 0.374 0.400 0.505 0.475
192 0.310 0.357 0.335 0.365 0.332 0.369 0.374 0.387 0.426 0.441 0.459 0.444 0.408 0.410 0.400 0.407 0.553 0.496
336 0.351 0.382 0.369 0.386 0.366 0.392 0.410 0.411 0.445 0.459 0.495 0.464 0.435 0.428 0.438 0.438 0.621 0.537
720 0.380 0.409 0.425 0.421 0.416 0.420 0.478 0.450 0.543 0.490 0.585 0.516 0.499 0.462 0.527 0.502 0.671 0.561
avg 0.328 0.371 0.357 0.378 0.351 0.380 0.400 0.406 0.448 0.452 0.481 0.456 0.429 0.425 0.435 0.437 0.588 0.517
ETTm2 96 0.163 0.254 0.167 0.269 0.165 0.255 0.187 0.267 0.203 0.287 0.192 0.274 0.189 0.280 0.209 0.308 0.255 0.339
192 0.218 0.290 0.224 0.303 0.220 0.292 0.249 0.309 0.269 0.328 0.280 0.339 0.253 0.319 0.311 0.382 0.281 0.340
336 0.271 0.328 0.281 0.342 0.274 0.329 0.321 0.351 0.325 0.366 0.334 0.361 0.314 0.357 0.442 0.466 0.339 0.372
720 0.356 0.382 0.397 0.421 0.362 0.385 0.408 0.403 0.421 0.415 0.417 0.413 0.414 0.413 0.675 0.587 0.433 0.432
avg 0.252 0.316 0.267 0.333 0.255 0.315 0.291 0.333 0.305 0.349 0.371 0.306 0.347 0.293 0.342 0.409 0.436 0.327
Weather 96 0.146 0.196 0.176 0.237 0.149 0.198 0.172 0.220 0.217 0.296 0.173 0.223 0.197 0.281 0.182 0.242 0.266 0.336
192 0.193 0.236 0.220 0.282 0.194 0.241 0.219 0.261 0.276 0.336 0.245 0.285 0.237 0.312 0.227 0.287 0.307 0.367
336 0.239 0.240 0.265 0.319 0.245 0.282 0.280 0.306 0.339 0.380 0.321 0.338 0.298 0.353 0.282 0.334 0.359 0.395
720 0.301 0.316 0.333 0.362 0.314 0.334 0.365 0.359 0.403 0.428 0.414 0.410 0.352 0.288 0.352 0.386 0.419 0.428
avg 0.219 0.247 0.248 0.300 0.225 0.264 0.259 0.287 0.309 0.360 0.288 0.314 0.271 0.334 0.261 0.312 0.338 0.382
Electricity 96 0.130 0.219 0.140 0.237 0.129 0.222 0.168 0.272 0.193 0.308 0.169 0.273 0.187 0.304 0.207 0.307 0.201 0.317
192 0.154 0.243 0.153 0.249 0.157 0.240 0.184 0.289 0.201 0.315 0.182 0.286 0.199 0.315 0.213 0.316 0.222 0.334
336 0.158 0.245 0.169 0.267 0.163 0.259 0.198 0.300 0.214 0.329 0.200 0.304 0.212 0.329 0.230 0.333 0.231 0.338
720 0.190 0.294 0.203 0.301 0.197 0.290 0.220 0.320 0.246 0.355 0.222 0.321 0.233 0.345 0.265 0.360 0.254 0.361
avg 0.158 0.251 0.166 0.263 0.161 0.252 0.192 0.295 0.214 0.327 0.193 0.296 0.208 0.323 0.229 0.329 0.227 0.338
Traffic 96 0.361 0.248 0.410 0.282 0.360 0.249 0.593 0.321 0.587 0.366 0.612 0.338 0.607 0.392 0.615 0.391 0.613 0.388
192 0.376 0.250 0.423 0.287 0.379 0.256 0.617 0.336 0.604 0.373 0.613 0.340 0.621 0.399 0.601 0.382 0.616 0.382
336 0.382 0.268 0.436 0.296 0.392 0.264 0.629 0.336 0.621 0.383 0.618 0.328 0.622 0.396 0.613 0.386 0.622 0.337
720 0.426 0.382 0.466 0.315 0.432 0.286 0.640 0.350 0.626 0.382 0.653 0.355 0.632 0.396 0.658 0.407 0.660 0.408
avg 0.386 0.261 0.433 0.295 0.390 0.263 0.620 0.336 0.610 0.376 0.628 0.379 0.624 0.340 0.621 0.396 0.622 0.392
ILI 24 1.301 0.739 2.215 1.081 1.319 0.754 2.317 0.934 3.228 1.260 2.294 0.945 2.527 1.020 8.313 2.144 3.483 1.287
36 1.409 0.824 1.963 0.963 1.430 0.834 1.972 0.920 2.679 1.080 1.825 0.848 2.615 1.007 6.631 1.902 3.103 1.148
48 1.520 0.807 2.130 1.024 1.553 0.815 2.238 0.940 2.622 1.078 2.010 0.900 2.359 0.972 7.299 1.982 2.669 1.085
60 1.553 0.852 2.368 1.096 1.520 0.788 2.027 0.928 2.857 1.157 2.178 0.963 2.487 1.016 7.283 1.985 2.770 1.125
avg 1.439 0.805 2.169 1.041 1.443 0.797 2.139 0.931 2.847 1.144 2.077 0.914 2.497 1.004 7.382 2.003 3.006 1.161
Table 12: Multivariate long-term time-series forecasting: lower values indicate better performance. Forecasting horizons are $h\in\{24,36,48,60\}$ for ILI and $h\in\{96,192,336,720\}$ for the rest.
Appendix G More Ablation Analysis
G.1 Look-Back Window Size vs Loss
In Figure 5, we have proved that RT works for the short look-back window better than the Dlinear method, since RT can capture the long-term dependencies using group reservoirs.
Figure 5: Look back window size vs loss
G.2 Graphical Explanation of Features
In Figure 6, we present a visual depiction that offers a comprehensive insight into the interpretability of features derived from both the reservoir output $\mathbf{z}_{t}$ and input feature $\mathbf{u}_{t}$.
G.3 Non-Linear output
In Table 13, we present a comparison of the non-linear readout of the reservoir using different activation functions.
Activation ETTh1 ETTh2 Traffic
Linear ReLu Tanh Self-Attention Linear ReLu Tanh Self-Attention Linear ReLu Tanh Self-Attention
MSE 0.538 0.462 0.474 0.441 0.447 0.608 0.495 0.485 0.498 0.428 0.463 0.426
MAE 0.562 0.470 0.475 0.455 0.348 0.552 0.468 0.478 0.327 0.293 0.314 0.281
Table 13: Comparison of MSE for different readout activation.
Appendix H Limitations
Our method assumes that the compressed memory represents the long context within the time-series data. However, in practice memory compression can result in potential information loss. Therefore, giving the whole context as input instead of compressing it to a fixed size may result in better performance. Additionally, we evaluated our method on standard time-series benchmarks and thus set the maximum horizon value as 720. There may exist datasets that have extremely long-term dependencies. To capture this dependency in such datasets, we would either have to increase the size of the reservoir’s input.